A sample of gas has a volume of 2.0 L at a pressure of 1.5 bar. If the gas is compressed to a volume of 1.0 L at constant temperature the pressure will be:


1
Expert's answer
2021-10-05T23:58:44-0400

"P_1V_1=P_2V_2"

"1.5bar \times 2.0L=P_2\times 1.0L"

"P_2= \frac{1.5bar \times 2.0L}{1.0L}=3.0bar"